Oc

Solve this second order Differential Equation using any method.
Answer #1

__0+.c!--10 01+ Cool +90 20 which is quadratic equation in a 0:-c1 172 4x914 2. 0,3-C-16= 4x914 -0,= -c+ V¢?4x314 öö caser.The solution is --- [=cfvalt 0 Cet [-(-VK) * aze Case II J 1f .c?.4420 - ec²-ug Lo Suppose w = ec² ug ie wio: 0 = -c IV-co (w